the power of a 2-mw laser beam decreases to 15 Mw after the beam traverses through 25km of a signal-mode optical fiber.calculate the attenuation the fiber
Answer #1

Solution:-

Given:
P1 = 2 mW = 2*10^-3
P2 = 15 \mu W = 15*10^-6
L = 25 km

Attenuation of the fibre = (10*log(P1/P2))L Attenuation of the fibre = (10*log ( 2*10^-3 / 15*10^-6)) / 25
Attenuation of the fibre = (10*log (2000/15)) / 25
Attenuation of the fibre = 10*log(133.333) / 25
Attenuation of the fibre = 0.8499 = 0.85 dB/km approximately